1. A body covers half a circle, of radius 1 m.  What is the ratio of distance to displacement of the body?
1
Expert's answer
2021-05-26T15:32:42-0400

Distance "s=2\pi r/2=\pi r"


Displacement "l=2r"


"s/l=\pi r/(2r)=\pi/2=3.14/2=1.57" . Answer